Multiply (4-7i)(5+6i)
Will it equal 62+11i, -42^2-11i+20, 62-11i, or -22-59i?

Answer by sheldonbbtrocks(115)

(4-7i)(5+6i)
20+24i-35i-42i^2
remember i^2=-1
-42(-1)=42
24i+20-35i+42
62-11i